OpenStudy (anonymous):

A special lottery is to be held to select the student who will live in the only deluxe room in a dormitory. There are seniors, juniors, and sophomores who applied. Each senior's name is placed in the lottery times; each junior's name, times; and each sophomore's name, time. What is the probability that a senior's name will be chosen?

OpenStudy (anonymous):

missing info

OpenStudy (anonymous):

A special lottery is to be held to select the student who will live in the only deluxe room in a dormitory. There are 100 seniors, 150 juniors, and 200 sophomores who applied. Each senior's name is placed in the lottery3 times; each junior's name,2 times; and each sophomore's name, 1 time. What is the probability that a senior's name will be chosen?

OpenStudy (anonymous):

hey everyone i got an answer.....:)

OpenStudy (anonymous):

well, we have 100 senior and each one got 3 tickets: 100*3=300 we have 150 juniors and each one got 2 tickets: 150*2=300 we have 200 sophmores and each one got 1 ticket: 200*1=200 This means we have 800 in total, and 300 seniors: making our probablilty 300/800 or 3/8. You just got schooled son